# EEG2Rep: Enhancing Self-supervised EEG Representation Through Informative Masked Inputs

### v1

@@ Abstract.
EEG Representation Learning, EEG Masking, EEG Classification
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Applied computing Health care information systemsccs: Computing methodologies Machine learning
@@ 1. Introduction
An electroencephalogram (EEG) is a noninvasive method that captures brain data by placing electrodes on the patient’s scalp surface, enabling the recording of electrical activity within the brain (EEG2002, ). This specialized and complex biological electrical signal serves as a reflection of the brain’s functional state, providing insights into the individual’s mental condition (EEG2005, ). From this data, valuable information can be extracted, including vital signs that facilitate continuous monitoring of the patient’s health (lotte2018review, ). Additionally, EEG plays a crucial role in diagnosing and identifying various brain conditions, finding applications in diverse healthcare domains such as sleep medicine, neurological disorders, cardiovascular disease detection, and activity monitoring (roy2019deep, ; chen2022brainnet, ).
@@
In the past decade, the integration of deep learning into biomedical research has experienced significant growth, demonstrating its ability to frequently outperform conventional machine learning methods across various tasks (roy2019deep, ; craik2019deep, ; hosseini2020review, ; weng2024self, ). However, the training of deep learning models for biomedical applications requires substantial amounts of data, annotated by experts, whose collection is often time and cost-prohibitive. Furthermore, deeper neural networks are susceptible to overfitting the EEG data, particularly in the presence of inter-subject variability (Bendr, ; schirrmeister2017deep, ). Self-supervised learning (SSL) has emerged as a prominent solution for such problems, as it allows learning powerful representations from vast unlabeled data by producing supervisory signals directly from the data (Bendr, ; maeeg, ; yang2023biot, ; weng2024self, ).
@@
In EEG analysis, two commonly used approaches for self-supervised learning are invariance-based and self-prediction methods (foumani2023deep, ; weng2024self, ). In invariance-based methods, the objective is to optimize an encoder to generate similar embeddings for two or more views of the same EEG time series (foumani2023deep, ). These different views of the time series are usually crafted using a set of manual data augmentations, including techniques like jittering, permutation, and scaling (TS-TCC2021, ; MohsenvandIM20, ). The main idea behind self-prediction methods is to remove or corrupt parts of the input and train the model to predict or reconstruct the altered content (foumani2023deep, ). For example, approaches like Masked AutoEncoders (MAE) (mae2022, ) learn representations by reconstructing randomly masked patches from an input (Bendr, ; mae2022, ; TST2021, ).
@@
Invariance-based pretraining methods can construct representations of high-level semantics by capturing essential features consistent across various data views. This strategy enables the model to identify and prioritize features important for understanding the underlying semantics of the data, as these features remain unchanged despite variations in the input. However, we believe that invariance-based methods perform well in computer vision and natural language processing due to the strong constraints present in image and text data. For example, the success with images arises from tasks related to object interpretation, where transformations such as scaling, blurring, and rotation assume that the resulting images will be similar to those generated in the original scenario with changes in camera zoom, stability, focus, or angle. However, there do not appear to be equivalent transformations that can be applied to EEG data. Augmentation methods applied to EEG data can inadvertently modify the semantic meaning, underlying distribution, and class representation of the signals (ling2022staging, ). For example, augmentation may introduce synthetic patterns or artifacts that do not align with genuine brain activity.
@@
To address the above-mentioned challenges we introduce EEG2Rep to enhance the semantic quality of EEG self-supervised representations without relying on prior knowledge about down-stream tasks. In contrast to the existing self-prediction methods that learn EEG representations by reconstructing the raw EEG data space (maeeg, ; Bendr, ), EEG2Rep is trained to reconstruct more abstract features of EEG data in the latent space. Such approach has shown to be effective in image and text representation learning (data2vec2-2023, ; i-jepa23, ; zhou2022image, ). Our motivation is that the existing noise in EEG is less likely to remain in the abstract features of EEG, and by learning to reconstruct the abstract features we potentially eliminate the unnecessary noise that exists in raw EEG data (addressing challenge 1). Additionally, as the abstract features are normalized within the representation space, the reconstruction of these features becomes more straightforward compared to reconstructing the potentially high amplitude value range of raw EEG data (addressing challenge 2). Finally, EEG2Rep leverages our novel Semantic Subsequence Preserving method to ensure that the context has sufficiently meaningful and rich semantic information (addressing challenge 3).
Fig. 1b displays the visualization of EEG representations learned by EEG2Rep. The two classes in this figure are easily separable which highlights the effectiveness of EEG2Rep in enhancing EEG representations and, consequently, improving classification accuracy. Another core component of EEG2Rep is its efficient multi-masking design. Specifically, we reuse the same target representation for various masked versions of each sample. Additionally, we predict the representation of various target blocks for a single masked input to further improve efficiency.
@@ 2.1. Problem Definition
Our goal is to address the problem of learning a nonlinear embedding function that can effectively map each EEG sample $X_{i}=\{x_{1},x_{2},\ldots,x_{L}\}$ from a given dataset $D$ into a concise and meaningful representation $R_{i}\in\mathbb{R}^{d_{e}}$, where $d_{e}$ indicates the desired representation dimension. The EEG dataset $D$ consists of $n$ samples, denoted as $D=\{{X_{1}},{X_{2}},...,{X_{n}}\}$, with each ${X_{i}}$ representing a $C$-channel EEG sequence of length $L$. To evaluate the quality of our learned representation $\mathbf{R}=\{R_{1},R_{2},\ldots,R_{n}\}$, we examine two scenarios based on the availability of labeled data: i) Linear Probing: We first pre-train a model without labels through a self-supervised pretext task. Upon completing the pre-training phase, we freeze the encoder and add a linear classifier on top of the pre-trained model’s output or intermediate representations. The linear classifier is then trained on a down-stream task, typically a classification task, utilizing the pre-trained representations as inputs, and ii) Fine-Tuning: Initially, we pre-train a model without labels through a self-supervised pretext task. Next, we perform fine-tuning by training the entire model for a few epochs using a labeled dataset in a fully supervised manner.
@@ 2.2. Model architecture
Building upon established works (schirrmeister2017deep, ; chambon2018deep, ; TS-TCC2021, ; ConvTran2023, ; series2vec23, ), we adopt a 3-layer convolutional neural network as input embedding to convert the raw EEG data into patches. Specifically, we feed EEG sample $X_{i}$ to the first layer that incorporates a depthwise convolutional layer, specifically designed to capture the spatial correlations between channels (schirrmeister2017deep, ; foumani2021disjoint, ). This is succeeded by a linear spatial filter to amplify the signal-to-noise ratio. The spatial filters leverage the fact that neural signals exhibit specific spatial patterns across the scalp, while noise sources may manifest more random spatial patterns (chambon2018deep, ).
Following the spatial filter, we integrate max pooling and spatial padding to ensure translation equivalence and effectively address edge effects. The output of this network is a set of EEG patches $\hat{S}_{\textbf{x}}=\{\hat{S}_{x_{1}},...,\hat{S}_{x_{l}}\}$ where $\hat{S}_{x_{i}}\in\mathbb{R}^{d_{x}\times l}$, and $d_{x}$ is the embedding dimension, and $l$ is the number of patches. Lastly, for every patch $\hat{S}_{x_{i}}$, the positional embedding feature of the $i^{th}$ position is added to it, resulting in the EEG patches $S_{\textbf{x}}=\{S_{x_{1}},...,S_{x_{l}}\}$. Please note for each EEG sequence $X_{i}$ we will have a set of EEG patches $S_{\textbf{x}}^{i}$ as the output of the input embedding network. However, for simplicity, we will drop the superscript $i$ from $S_{\textbf{x}}^{i}$ and we use $S_{\textbf{x}}$ as an input to the following components in the EEG2Rep architecture.
@@
Previous self-prediction methods for EEG data have primarily followed the approach of Masked Autoencoders (MAE) (mae2022, ), reconstructing local windows of the raw input EEG (maeeg, ), or adopted a BERT-like method (bert2018, ), predicting discrete representations (Bendr, ). However, the resulting representations often demonstrate lower semantic quality than invariance-based methods during off-the-shelf evaluations, such as linear probing, due to the intrinsic characteristics of EEG data mentioned earlier. Low signal-to-noise ratio in EEG challenges the reconstruction task and a wide range of amplitudes in EEG further complicates the optimization process.
@@
EEG2Rep comprises three main components: Target network uses the complete (unmasked) input embedding $S_{\textnormal{x}}$ to generate semantically rich representations, allowing each patch to encode knowledge of all others through its self-attention architecture. The resulting output serves as the target for our learning task. Context network shares the architecture with the target network, differing only in its use of the masked version of the input embedding to generate representations for visible patches. Our innovative semantic subsequence preserving method ensures that the context network’s output contains sufficient semantic information for reconstruction purposes. We use a standard transformer (attention2017, ; bert2018, ) for both target and context networks. Predictor network leverages the output of the context network and randomly chosen masked tokens to regress the targets, i.e., the output of the target network. We enhance contextualized information integration for the predictor network by incorporating cross-attention-based transformers (bert2018, ). Below, we elaborate on the process of creating the target, context, and predictor networks for the training task, providing clarification on how these networks are parameterized.
@@
Given the output of the context encoder $\textbf{r}(q)$, and a target mask $B_{i}$, we wish to predict $M$ target block representations $\{\textbf{y}(1),\ldots,\textbf{y}(M)\}$. To that end, for a given target block $\textbf{y}(i)$, the predictor $g_{\phi}$ takes as input the output of the context encoder $\textbf{r}(q)$ and a mask token $m(i)$ for each patch we wish to predict and outputs a patch-level prediction:
@@ 2.3. Semantic Subsequence Preserving (SSP)
Figure 3. Different masking strategies applied to EEG samples with a 50% masking ratio. From top to bottom: (1) Original EEG sample, (2) Random masking, (3) Block masking, and (4) Semantic Subsequence Preserving (SSP).
Random masking has proven successful for Masked Autoencoders (mae2022, ), where random patches are sampled without replacement, following a uniform distribution, resulting in significant efficiency improvements. Fig. 3 illustrates various masking strategies applied to EEG samples, with the top subplot showcasing the original EEG sample recorded for 0.5 seconds at 128 Hz. The subsequent subplots demonstrate different masking techniques, each with a $50\%$ masking ratio. The second subplot exhibits random masking, following the Masked Autoencoders style. While the MAE approach has demonstrated success in computer vision tasks, it may pose challenges in building semantic representations for EEG data due to the lack of structure in the created masks (e.g., random binary masking). Notably, most of the masked regions can be reconstructed trivially, for example, through interpolation. On the other hand, block masking (data2vec2022, ; i-jepa23, ), involves masking entire blocks of time-steps or patches. However, this approach does not guarantee that large continuous portions of the training sample remain unmasked. As shown in Fig. 3, block masking primarily focuses on the structure of the mask rather than context. Our objective is to enable the context encoder to construct semantically rich representations over local regions of the sample.
@@
Here, $l$ represents the total number of time steps (patches) in a training sample, $\beta$ is the desired number of blocks, and $\rho$ is the mask ratio. We allow visible blocks to overlap, leading to over size subsequences preserving and some variability in the number of visible time steps for each EEG sample. As shown in Fig. 3, the first preserved subsequence arises from overlapping two blocks. As we encode only visible time steps, we adopt a simple strategy to maintain uniformity in the number of masked time steps across all samples in a batch. Following the semantic subsequence preserving step, we randomly preserve individual time steps until we reach the targeted number of preserved time steps, calculated as $(1-\rho)\times l$.
@@ 2.4. Efficiency: Multiple masking
Multiple Prediction Subsequently, having obtained representations of various masked versions of the same input, we predict the representation of various target blocks for a single masked input to further enhance efficiency.
Now, the reconstruction loss in Equation 4 is updated to:
@@ 3. Related Work
Traditional approaches to extract useful features from EEG are surveyed in (subha2010eeg, ). In recent years, advanced self-supervised learning methods have been proposed, aiming to derive representations from sparsely labeled EEG data. Here, we discuss two primary self-supervised approaches for EEG representation learning: invariance-based and self-prediction, and refer interested readers to (weng2024self, ; foumani2023deep, ) for more details.
@@ 3.2. Self-Prediction based self-supervised learning
Emotiv
DREAMER (dreamer, ) 128Hz 14 2s 77,910 Emotion Detection
@@
Temple
TUAB (TUAB, ) 256Hz 16 10s 409,455 Abnormal EEG Classification

### v2

@@ Abstract.
EEG Representation Learning, EEG self-supervised Learning, EEG Masking, EEG Classification
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671600isbn: 979-8-4007-0490-1/24/08submissionid: ads0493ccs: Applied computing Health care information systemsccs: Computing methodologies Machine learning
@@ 1. Introduction
An electroencephalogram (EEG) is a noninvasive method that captures brain data by placing electrodes on the patient’s scalp surface, enabling the recording of electrical activity within the brain (EEG2002, ). This specialized and complex biological electrical signal serves as a reflection of the brain’s functional state, providing insights into the individual’s mental condition (EEG2005, ). From this data, valuable information can be extracted, including vital signs that facilitate continuous monitoring of the patient’s health (lotte2018review, ). Additionally, EEG plays a crucial role in diagnosing and identifying various brain conditions, finding applications in diverse healthcare domains such as sleep medicine, neurological disorders, cardiovascular disease detection, and activity monitoring (roy2019deep, ; chen2022brainnet, ; rakovic2024measuring, ).
@@
In the past decade, the integration of deep learning into biomedical research has experienced significant growth, demonstrating its ability to frequently outperform conventional machine learning methods across various tasks (roy2019deep, ; craik2019deep, ; hosseini2020review, ; rakovic2024measuring, ; weng2024self, ). However, the training of deep learning models for biomedical applications requires substantial amounts of data, annotated by experts, whose collection is often time and cost-prohibitive. Furthermore, deeper neural networks are susceptible to overfitting the EEG data, particularly in the presence of inter-subject variability (Bendr, ; schirrmeister2017deep, ). Self-supervised learning (SSL) has emerged as a prominent solution for such problems, as it allows learning powerful representations from vast unlabeled data by producing supervisory signals directly from the data (Bendr, ; maeeg, ; yang2023biot, ; weng2024self, ).
@@
In EEG analysis, two commonly used self-supervised learning approaches are invariance-based methods and self-prediction methods (foumani2023deep, ; weng2024self, ). In invariance-based methods, the objective is to optimize an encoder to generate similar embeddings for two or more views of the same EEG time series (foumani2023deep, ). These different views of the time series are usually crafted using a set of manual data augmentations, including techniques like jittering, permutation, and scaling (TS-TCC2021, ; MohsenvandIM20, ). The main idea behind self-prediction methods is to remove or corrupt parts of the input and train the model to predict or reconstruct the altered content (foumani2023deep, ). For example, approaches like Masked AutoEncoders (MAE) (mae2022, ) learn representations by reconstructing randomly masked patches from an input (Bendr, ; mae2022, ; TST2021, ).
@@
Invariance-based pretraining methods can construct representations of high-level semantics by capturing essential features consistent across various data views. This strategy enables the model to identify and prioritize features important for understanding the underlying semantics of the data, as these features remain unchanged despite variations in the input. However, we believe that invariance-based methods perform well in computer vision and natural language processing due to the strong constraints present in image and text data. For example, the success with images arises from tasks related to object interpretation, where transformations such as scaling, blurring, and rotation assume that the resulting images will be similar to those generated in the original scenario with changes in camera zoom, stability, focus, or angle. However, there do not appear to be equivalent transformations that can be applied to EEG data. Augmentation methods applied to EEG data can inadvertently modify the semantic meaning, underlying distribution, and class representation of the signals (ling2022staging, ; series2vec23, ). For example, augmentation may introduce synthetic patterns or artifacts that do not align with genuine brain activity.
@@
To address the challenges mentioned above, we introduce EEG2Rep to enhance the semantic quality of EEG representations without relying on prior knowledge about downstream tasks. In contrast to the existing self-prediction methods that learn EEG representations by reconstructing the raw EEG data space (maeeg, ; Bendr, ), EEG2Rep is trained to reconstruct more abstract features of EEG data in the latent space. Such approaches have shown to be effective in image and text representation learning (data2vec2-2023, ; i-jepa23, ; zhou2022image, ). Our motivation is that the existing noise in EEG is less likely to remain in the abstract features of EEG, and by learning to reconstruct the abstract features we potentially eliminate the unnecessary noise that exists in raw EEG data (addressing challenge 1). Additionally, as the abstract features are normalized within the representation space, the reconstruction of these features becomes more straightforward compared to reconstructing the potentially high amplitude value range of raw EEG data (addressing challenge 2). Finally, EEG2Rep leverages our novel Semantic Subsequence Preserving method to ensure that the context has sufficiently meaningful and rich semantic information (addressing challenge 3).
Fig. 1b displays the visualization of EEG representations learned by EEG2Rep. The two classes in this figure are easily separable which highlights the effectiveness of EEG2Rep in enhancing EEG representations and, consequently, improving classification accuracy. Another core component of EEG2Rep is its efficient multi-masking design. Specifically, we reuse the same target representation for various masked versions of each sample. Additionally, we predict the representation of various target blocks for a single masked input to improve efficiency further.
@@ 2.1. Problem Definition
Our goal is to address the problem of learning a nonlinear embedding function that can effectively map each EEG sample $X_{i}=\{x_{1},x_{2},\ldots,x_{L}\}$ from a given dataset $D$ into a concise and meaningful representation $R_{i}\in\mathbb{R}^{d_{e}}$, where $d_{e}$ indicates the desired representation dimension. The EEG dataset $D$ consists of $n$ samples, denoted as $D=\{{X_{1}},{X_{2}},...,{X_{n}}\}$, with each ${X_{i}}$ representing a $C$-channel EEG sequence of length $L$. To evaluate the quality of our learned representation $\mathbf{R}=\{R_{1},R_{2},\ldots,R_{n}\}$, we examine two scenarios based on the availability of labeled data: i) Linear Probing: We first pre-train a model without labels through a self-supervised pretext task. Upon completing the pre-training phase, we freeze the encoder and add a linear classifier on top of the pre-trained model’s output or intermediate representations. The linear classifier is then trained on a downstream task, typically a classification task, utilizing the pre-trained representations as inputs, and ii) Fine-Tuning: Initially, we pre-train a model without labels through a self-supervised pretext task. Next, we perform fine-tuning by training the entire model for a few epochs using a labeled dataset in a fully supervised manner.
@@ 2.2. Model architecture
Building upon established works (schirrmeister2017deep, ; chambon2018deep, ; foumani2021disjoint, ; TS-TCC2021, ; ConvTran2023, ; series2vec23, ), we adopt a 3-layer convolutional neural network as input embedding to convert the raw EEG data into patches. Specifically, we feed EEG sample $X_{i}$ to the first layer that incorporates a depthwise convolutional layer, specifically designed to capture the spatial correlations between channels (schirrmeister2017deep, ; foumani2021disjoint, ). This is succeeded by a linear spatial filter to amplify the signal-to-noise ratio. The spatial filters leverage the fact that neural signals exhibit specific spatial patterns across the scalp, while noise sources may manifest more random spatial patterns (chambon2018deep, ).
Following the spatial filter, we integrate max pooling and spatial padding to ensure translation equivalence and effectively address edge effects. The output of this network is a set of EEG patches $\hat{S}_{\textbf{x}}=\{\hat{S}_{x_{1}},...,\hat{S}_{x_{l}}\}$ where $\hat{S}_{x_{i}}\in\mathbb{R}^{d_{x}}$, and $d_{x}$ is the embedding dimension, and $l$ is the number of patches. Lastly, for every patch $\hat{S}_{x_{i}}$, the positional embedding feature of the $i^{th}$ position is added to it, resulting in the EEG patches $S_{\textbf{x}}=\{S_{x_{1}},...,S_{x_{l}}\}$ (shown in blue in Fig. 2). Please note for each EEG sequence $X_{i}$, we will have a set of EEG patches $S_{\textbf{x}}^{i}$ as the output of the input embedding network. However, for simplicity, we will drop the superscript $i$ from $S_{\textbf{x}}^{i}$ and use $S_{\textbf{x}}$ as an input to the following components in the EEG2Rep architecture.
@@
Previous self-prediction methods for EEG data have primarily followed the approach of Masked Autoencoders (MAE) (mae2022, ), reconstructing local windows of the raw input EEG (maeeg, ), or have adopted a BERT-like method (bert2018, ), predicting discrete representations (Bendr, ). However, the resulting representations often demonstrate lower semantic quality than invariance-based methods during off-the-shelf evaluations, such as linear probing, due to the intrinsic characteristics of EEG data mentioned earlier. Low signal-to-noise ratio in EEG challenges the reconstruction task and a wide range of amplitudes in EEG further complicates the optimization process.
@@
EEG2Rep comprises three main components: Target network uses the complete (unmasked) input embedding $S_{\textbf{x}}$ to generate semantically rich representations, allowing each patch to encode knowledge of all others through its self-attention architecture. The resulting output serves as the target for our learning task. Context network shares the architecture with the target network, differing only in its use of the masked version of the input embedding to generate representations for visible patches. Our innovative semantic subsequence preserving method ensures that the context network’s output contains sufficient semantic information for reconstruction purposes. We use a standard transformer (attention2017, ; bert2018, ) for both target and context networks. Predictor network leverages the output of the context network and randomly chosen masked tokens to regress the targets, i.e., the output of the target network. We enhance contextualized information integration for the predictor network by incorporating cross-attention-based transformers (bert2018, ). Below, we elaborate on the process of creating the target, context, and predictor networks for the training task, providing clarification on how these networks are parameterized.
@@
Given the output of the context encoder $\textbf{r}(q)$, our goal is to predict $M$ target block representations $\{\textbf{y}(1),\ldots,\textbf{y}(M)\}$. For each target block $\textbf{y}(i)$, the predictor $g_{\phi}$ takes as input the output of the context encoder $\textbf{r}(q)$ and a mask token $m(i)$ for each patch we wish to predict and outputs a patch-level prediction:
@@ 2.3. Semantic Subsequence Preserving (SSP)
Figure 3. Different masking strategies applied to EEG samples from Crowdsourced datasets (crowdsourced, ) with a 50% masking ratio. From top to bottom: (1) Original EEG sample, (2) Random masking, (3) Block masking, and (4) Semantic Subsequence Preserving (SSP).
Random masking has proven successful for Masked Autoencoders (mae2022, ), where random patches are sampled without replacement, following a uniform distribution, resulting in significant efficiency improvements. Fig. 3 illustrates various masking strategies applied to EEG samples, with the top subplot showcasing the original EEG sample from Crowdsourced datasets (crowdsourced, ) recorded for 0.5 seconds at 128 Hz. The subsequent subplots demonstrate different masking techniques, each with a $50\%$ masking ratio. The second subplot exhibits random masking, following the Masked Autoencoders style. While the MAE approach has demonstrated success in computer vision tasks, it may pose challenges in building semantic representations for EEG data due to the lack of structure in the created masks (e.g., random binary masking). Notably, most of the masked regions can be reconstructed trivially, for example, through interpolation. On the other hand, block masking (data2vec2022, ; i-jepa23, ), involves masking entire blocks of time-steps or patches. However, this approach does not guarantee that large continuous portions of the training sample remain unmasked. As shown in Fig. 3, block masking primarily focuses on the structure of the mask rather than context. Our objective is to enable the context encoder to construct semantically rich representations over local regions of the sample.
@@
Here, $l$ represents the total number of time steps (patches) in a training sample, $\beta$ is the desired number of blocks, and $\rho$ is the mask ratio. We allow visible blocks to overlap, leading to oversized subsequences preserving and some variability in the number of visible time steps for each EEG sample. As shown in Fig. 3, the first preserved subsequence arises from overlapping two blocks. As we encode only visible time steps, we adopt a simple strategy to maintain uniformity in the number of masked time steps across all samples in a batch. Following the semantic subsequence preserving step, we randomly preserve individual time steps until we reach the targeted number of preserved time steps, calculated as $(1-\rho)\times l$.
@@ 2.4. Efficiency: Multiple masking
Multiple Prediction Subsequently, having obtained representations of various masked versions of the same input, we predict the representation of various target blocks for a single masked input to further enhance efficiency. Now, the reconstruction loss in Equation 4 is updated to:
@@ 3. Related Work
Traditional approaches to extract useful features from EEG are surveyed in (subha2010eeg, ). In recent years, advanced self-supervised learning methods have been proposed aiming to derive representations from sparsely labeled EEG data. Here, we discuss two primary self-supervised approaches for EEG representation learning: invariance-based and self-prediction, and refer interested readers to (weng2024self, ; foumani2023deep, ) for more details.
@@ 3.2. Self-Prediction based self-supervised learning
Emotiv DREAMER (dreamer, ) 128Hz 14 2s 77,910 Emotion Detection
@@
Temple TUAB (TUAB, ) 256Hz 16 10s 409,455 Abnormal EEG Classification